Differentiate y= In ( 2cot²x )


1
Expert's answer
2022-03-19T02:36:50-0400

"y'=\frac{1}{2cot^2(x)}\cdot 2\cdot2cot(x)\cdot (-\frac{1}{sin^2(x)})=\\
=-\frac{2}{sin^2(x)cot(x)}=-\frac{2}{sinxcosx}"